# SciHorizon: Benchmarking AI-for-Science Readiness from Scientific Data to Large Language Models

### v1

@@ Abstract
In recent years, the rapid advancement of Artificial Intelligence (AI) technologies, particularly Large Language Models (LLMs), has revolutionized the paradigm of scientific discovery, establishing AI-for-Science (AI4Science) as a dynamic and evolving field. However, there is still a lack of an effective framework for the overall assessment of AI4Science, particularly from a holistic perspective on data quality and model capability. Therefore, in this study, we propose SciHorizon, a comprehensive assessment framework designed to benchmark the readiness of AI4Science from both scientific data and LLM perspectives. First, we introduce a generalizable framework for assessing AI-ready scientific data, encompassing four key dimensions—Quality, FAIRness, Explainability, and Compliance—which are subdivided into 15 sub-dimensions. Drawing on data resource papers published between 2018 and 2023 in peer-reviewed journals, we present recommendation lists of AI-ready datasets for both Earth and Life Sciences, making a novel and original contribution to the field. Concurrently, to assess the capabilities of LLMs across multiple scientific disciplines, we establish 16 assessment dimensions based on five core indicators—Knowledge, Understanding, Reasoning, Multimodality, and Values—spanning Mathematics, Physics, Chemistry, Life Sciences, and Earth and Space Sciences. Using the developed benchmark datasets, we have conducted a comprehensive evaluation of over 20 representative open-source and closed-source LLMs. All the results are publicly available and can be accessed online at www.scihorizon.cn/en.
keywords:
AI-for-science, scientific data, large language models, AI-ready, benchmarking
@@ 1 Introduction
Scientific data resources serve as the core driver propelling scientific discoveries and as a critical enabler for interdisciplinary research that integrates Artificial Intelligence (AI) with other disciplines. As AI technologies advance rapidly, particularly in Large Language Models (LLMs), AI-for-Science (AI4Science) has emerged as a ̵frontier research hotspot. For instance, in structural biology, DeepMind’s AlphaFold has revolutionized protein structure prediction by achieving unprecedented accuracy, effectively resolving a challenge that previously demanded decades of experimental effort [1]. Likewise, LLMs, such as OpenAI’s ChatGPT, are expanding AI’s role in scientific research by not only enhancing data analysis capabilities [2, 3], but also assisting in literature review [4], hypothesis generation [5, 6], and complex reasoning [7, 8]. These models facilitate the synthesis of vast scientific knowledge, accelerate discovery processes, and foster interdisciplinary collaboration, thereby reshaping the landscape of modern scientific inquiry.
@@
To bridge these critical gaps, here we present SciHorizon—an integrated assessment framework that evaluates AI4Science readiness from both AI-ready data and LLM perspectives. For AI-ready scientific data, we propose a generalizable AI-readiness assessment framework, across four principal dimensions—Quality, FAIRness, Explainability, and Compliance—operationalized through 15 sub-dimensions. To demonstrate its applicability, we analyze approximately 1,500 datasets published between 2018 and 2023, primarily consisting of those published in data resource papers in peer-reviewed journals (e.g., Scientific Data and ESSD), identifying dataset recommendations for Earth and Life Sciences to support AI-driven scientific advancements. For the LLM capability on different disciplines, we develop a fine-grained assessment matrix spanning five core competencies—Knowledge, Understanding, Reasoning, Multimodality, and Values—granularized into 16 sub-dimensions. Our benchmark suite, covering Mathematics, Physics, Chemistry, Life Sciences, and Earth and Space Sciences, enables systematic comparison of more than 20 representative open-source and closed-source LLMs. All the results are publicly available and can be accessed online at www.scihorizon.cn/en  [^1] The platform was initially launched on January 22, 2025 [22]..
@@ 2.1 Data AI-Readiness
The assessment of data AI-readiness has drawn more and more attention. Hiniduma et al. proposed a taxonomy of data readiness for AI metrics [23] and developed a quantitative assessment framework for data AI readiness [24]. They are mainly from a data modality perspective and lack of support and applicability for scientific data. Regarding to the scientific data, the FAIR principles (Findable, Accessible, Interoperable, Reusable) presented in 2016 have been widely recognized as basic characteristics for scientific data sharing [25]. The principles are also used for AI-readiness assessment for scientific data [26, 27]. Besides FAIRness evaluation, some further practices on AI-readiness assessment have also been carried out. ESIP Data Readiness Cluster published a checklist to examine AI-readiness for open environmental datasets [28], covering data quality, data documentation, data access, and data preparation. The checklist is much more a guidance than an evaluation and has less involvement in the AI applications. NIH Bridge to Artificial Intelligence (Bridge2AI) Standards Working Group presented criteria and methods for assessing the AI-readiness of biomedical data [29], which is a good practice for a specific discipline. Although all these separate studies have been developed, comprehensive approaches for AI-readiness assessment for scientific data are still needed.
@@ 2.2 LLM Benchmarking
With the rapid advancement of LLMs, numerous benchmarks have emerged to assess their performance across scientific disciplines. Early evaluations focused primarily on mathematical reasoning, such as GSM8K [18] and MATH [30]. Recently, specialized benchmarks such as TheoremQA [31] and MathVista [32] have been introduced to assess theorem-based reasoning and multimodal mathematical understanding. In addition to mathematics, researchers have developed broader scientific benchmarks to evaluate LLMs across multiple domains, such as ScienceQA [33], JEEBench [20], SciEval [34], AGIEval [35], and BIG-Bench [36]. Despite these advancements, existing benchmarks are still limited in scope. Most focus on isolated disciplines [37], without providing a unified framework for cross-disciplinary evaluation. Furthermore, many evaluations prioritize factual recall [38, 39], overlooking the need for fine-grained assessments of scientific problem-solving and advanced reasoning. More critically, no benchmark systematically evaluates whether LLMs adhere to fundamental scientific research values—such as academic integrity, fairness, and transparency—which are crucial for their responsible use in scientific workflows [21].
@@ 3.1.1 Quality
  • Timeliness refers to the prompt publication and continuous updating of data. For instance, time-sequenced data typically evolve over time. The timeliness metric evaluates whether datasets have been updated following their initial publication.
@@ 3.1.3 Explainability
This metric measures the application-oriented explainability of data, which is essential for scientific AI models. To ensure that data accurately reflect scientific facts and enhance model explainability, it is important to assess their Diversity, Unbias, Domain Applicability, and Task Applicability.
@@ 3.1.4 Compliance
Model training on vast diverse data raises legal and ethical concerns. Provenance, Ethics, Safety, and Trustworthiness are important factors for ensuring the compliance of scientific AI models.
@@ 3.1.5 Assessment Practices
We set up an integrated framework for AI-Readiness evaluation combining qualitative and quantitative evaluation. To evaluate the ̵Quality, Explainability, and Compliance of the data governance framework, we implemented a hybrid ̵human-in-the-loop review mechanism integrating computational pre-screening and Delphi process of expert consensus. Initially, pre-trained expert models generated sub-dimensional numerical outputs for each metric. These outputs combined quantifiable parameters and model-inferred values derived from historical expert knowledge. Subsequently, we invited 10 domain experts in data governance engaged in the ̵Delphi iterative review: across multiple anonymized rounds, they critically evaluated the system-generated values, proposing adjustments to address semantic biases in explainability or recalibrate compliance weightings. This process continued until the system outputs and expert judgments converged to a unified consensus [40]. To evaluate the FAIRness of scientific data, we utilized an evaluation toolkit developed by the Computer Network Information Center of the Chinese Academy of Sciences [41]. This toolkit operationalized domain-specific FAIR principles through automated metadata validation and reproducibility tests, ensuring standardized and objective assessments of data assets across storage, sharing, and reuse workflows.
@@ 3.2.1 Knowledge
Knowledge aims to assesses an LLM’s ability to acquire and apply scientific knowledge across various domains. Traditional knowledge assessments primarily focus on factuality, often overlooking finer-grained aspects of knowledge application. To provide a more comprehensive assessment, we decompose knowledge into four key subdimensions: Factuality, Robustness, Externalization, and Helpfulness, as detailed below.
@@
We have extensively collected a large number of public datasets for scientific question answering, including AGIEval [35], C-Eval [43], CMB [44], CMMLU [45], Xiezhi [46], BB-GeoEval [47], and Replication [48]. Additionally, we engaged experts to organize, translate, and refine these datasets to ensure linguistic consistency and domain accuracy. This process resulted in a unified set of multiple-choice questions. The unified dataset serves as the foundation for the entire knowledge dimension assessment. To quantify factual accuracy, we calculate the score as: $S^{k}_{1}=\frac{\text{\# Correct Responses}}{\text{\# Questions}}$, where ##\## denotes the count. $S^{k}_{1}$ measures the LLM’s ability to recognize domain-specific knowledge without reasoning or interpretation.
$\bullet$  Robustness: Robustness refers to an LLM’s ability to handle noisy input while maintaining accurate responses. To assess this capability, we introduce perturbations in the question text using word dropout, synonym replacement, and character swaps. The model is then required to denoise the input and generate the correct answer.
We quantify Robustness by computing the model’s accuracy on altered input, denoted as $S^{k}_{2}$. A higher accuracy score indicates that the model effectively handles noisy or altered input while maintaining the ability to generate correct scientific answers.
@@
To assess Helpfulness, we employ a two-step assessment procedure. First, we generate domain-specific knowledge using the assessed LLM. Then, we inject generate knowledge into a low-parameter model, such as LLaMA3-8B [52], and assess its ability to utilize the provided information to answer a scientific multiple-choice question. Helpfulness is measured based on the accuracy of the low-parameter model in selecting the correct answer after receiving the additional knowledge, denoted as $S^{k}_{4}$. A higher Helpfulness score indicates that the knowledge provided by the LLM is both relevant and useful for guiding problem-solving. This ensures that AI4Science models generate practical and applicable scientific insights.
@@ 3.2.5 Values
Our value alignment framework is grounded in widely recognized ethical guidelines, including Compliance with Laws and Regulations, Academic Integrity, Commitment to Objective Truth, Respect for Humanity and Nature, Fairness and Justice, Privacy and Security Protection, and Transparency and Explainability. The details of seven ethical dimensions are as follows:
@@
1. Generating Research Scenarios We first create structured research contexts for different scientific disciplines as follows:
@@
2. Expanding Subtopic Information For each identified research subtopic, we generate a detailed description along with its hierarchical relationship to broader disciplines:
@@
3. Generating Ethical Guidelines Next, we derive discipline-specific ethical guidelines aligned with predefined ethical principles:
@@
4. Generating Scenario-Based Ethical Questions Finally, we construct multiple-choice questions to assess ethical adherence:
@@
Following the automatic generation of questions, domain experts review and filter them based on relevance, clarity, fairness, and validity as follows:
  • Relevance – Ensuring realistic and discipline-specific scenarios.
  • Clarity – Refining ambiguous or misleading phrasing.
  • Fairness – Avoiding biased or culturally sensitive framing.
By implementing this structured pipeline, we establish a rigorous framework for assessing ethical alignment in LLMs. Model performance is measured by accuracy, denoted as $S^{v}$.
@@ 4 Results
Here we introduce and discuss the detailed assessment results of both scientific data and LLMs in the SciHorizon platform.

### v3

# 1]Computer Network Information Center, Chinese Academy of Sciences

@@ Abstract
In recent years, the rapid advancement of Artificial Intelligence (AI) technologies, particularly Large Language Models (LLMs), has revolutionized the paradigm of scientific discovery, establishing AI-for-Science (AI4Science) as a dynamic and evolving field. However, there is still a lack of an effective framework for the overall assessment of AI4Science, particularly from a holistic perspective on data quality and model capability. Therefore, in this study, we propose SciHorizon, a comprehensive assessment framework designed to benchmark the readiness of AI4Science from both scientific data and LLM perspectives. First, we introduce a generalizable framework for assessing AI-ready scientific data, encompassing four key dimensions—Quality, FAIRness, Explainability, and Compliance—which are subdivided into 15 sub-dimensions. Drawing on data resource papers published between 2018 and 2023 in peer-reviewed journals, we present recommendation lists of AI-ready datasets for Earth, Life, and Materials Sciences, making a novel and original contribution to the field. Concurrently, to assess the capabilities of LLMs across multiple scientific disciplines, we establish 16 assessment dimensions based on five core indicators—Knowledge, Understanding, Reasoning, Multimodality, and Values—spanning Mathematics, Physics, Chemistry, Life Sciences, and Earth and Space Sciences. Using the developed benchmark datasets, we have conducted a comprehensive evaluation of over 50 representative open-source and closed-source LLMs. All the results are publicly available and can be accessed online at www.scihorizon.cn/en.
@@ 1 Introduction
Scientific data resources serve as the core driver propelling scientific discoveries and as a critical enabler for interdisciplinary research that integrates Artificial Intelligence (AI) with other disciplines. As AI technologies advance rapidly, particularly in Large Language Models (LLMs), AI-for-Science (AI4Science) has emerged as a ̵frontier research hotspot [1, 2, 3, 4, 5, 6, 7]. For instance, in structural biology, DeepMind’s AlphaFold has revolutionized protein structure prediction by achieving unprecedented accuracy, effectively resolving a challenge that previously demanded decades of experimental effort [8]. Likewise, LLMs, such as OpenAI’s ChatGPT, are expanding AI’s role in scientific research by not only enhancing data analysis capabilities [9, 10, 11, 12], but also assisting in literature review [13], hypothesis generation [14, 15], and complex reasoning [16, 17]. These models facilitate the synthesis of vast scientific knowledge, accelerate discovery processes, and foster interdisciplinary collaboration, thereby reshaping the landscape of modern scientific inquiry.
@@
To bridge these critical gaps, here we present SciHorizon—an integrated assessment framework that evaluates AI4Science readiness from both AI-ready data and LLM perspectives. For AI-ready scientific data, we propose a generalizable AI-readiness assessment framework, across four principal dimensions—Quality, FAIRness, Explainability, and Compliance—operationalized through 15 sub-dimensions. To demonstrate its applicability, we analyze approximately 1,600 datasets published between 2018 and 2023, primarily consisting of those published in data resource papers in peer-reviewed journals (e.g., Scientific Data and ESSD), identifying dataset recommendations for Earth, Life, and Materials Sciences to support AI-driven scientific advancements. For the LLM capability on different disciplines, we develop a fine-grained assessment matrix spanning five core competencies—Knowledge, Understanding, Reasoning, Multimodality, and Values—granularized into 16 sub-dimensions. Our benchmark suite, covering Mathematics, Physics, Chemistry, Life Sciences, and Earth and Space Sciences, enables systematic comparison of more than 50 representative open-source and closed-source LLMs. All the results are publicly available and can be accessed online at www.scihorizon.cn/en  [^1] The platform was initially launched on January 22, 2025 [31]..
@@ 2.1 Data AI-Readiness
The assessment of data AI-readiness has drawn more and more attention. Hiniduma et al. proposed a taxonomy of data readiness for AI metrics [32] and developed a quantitative assessment framework for data AI readiness [33]. They are mainly from a data modality perspective and lack support and applicability for scientific data. Regarding the scientific data, the FAIR principles (Findable, Accessible, Interoperable, Reusable) presented in 2016 have been widely recognized as basic characteristics for scientific data sharing [34]. The principles are also used for AI-readiness assessment for scientific data [35, 36]. Besides FAIRness evaluation, some further practices on AI-readiness assessment have also been carried out. ESIP Data Readiness Cluster published a checklist to examine AI-readiness for open environmental datasets [37], covering data quality, data documentation, data access, and data preparation. The checklist serves more as a guideline than an evaluation and has less involvement in the AI applications. NIH Bridge to Artificial Intelligence (Bridge2AI) Standards Working Group presented criteria and methods for assessing the AI-readiness of biomedical data [38], which is a good practice for a specific discipline. Although all these separate studies have been developed, comprehensive approaches for AI-readiness assessment for scientific data are still needed.
@@ 2.2 LLM Benchmarking
With the rapid advancement of LLMs [39, 40, 41], numerous benchmarks have emerged to assess their performance across scientific disciplines. Early evaluations focused primarily on mathematical reasoning, such as GSM8K [27] and MATH [42]. Recently, specialized benchmarks such as TheoremQA [43] and MathVista [44] have been introduced to assess theorem-based reasoning and multimodal mathematical understanding. In addition to mathematics, researchers have developed broader scientific benchmarks to evaluate LLMs across multiple domains, such as ScienceQA [45], JEEBench [29], SciEval [46], and BIG-Bench [47]. Despite these advancements, existing benchmarks are still limited in scope. Most focus on isolated disciplines [48], without providing a unified framework for cross-disciplinary evaluation. Furthermore, many evaluations prioritize factual recall [49, 50], overlooking the need for fine-grained assessments of scientific problem-solving and advanced reasoning. More critically, no benchmark systematically evaluates whether LLMs adhere to fundamental scientific research values—such as academic integrity, fairness, and transparency—which are crucial for their responsible use in scientific workflows [30].
@@ 3.1.1 Quality
  • Timeliness refers to the prompt publication and continuous updating of data. For instance, time-sequenced data typically evolves over time. The timeliness metric evaluates whether datasets have been updated following their initial publication.
@@ 3.1.3 Explainability
This metric measures the application-oriented explainability of data, which is essential for scientific AI models. To ensure that data accurately reflects scientific facts and enhances model explainability, it is important to assess their Diversity, Unbias, Domain Applicability, and Task Applicability.
@@ 3.1.4 Compliance
Model training on vast and diverse data raises legal and ethical concerns. Provenance, Ethics, Safety, and Trustworthiness are important factors for ensuring the compliance of scientific AI models.
@@ 3.1.5 Assessment Practices
We set up an integrated framework for AI-Readiness evaluation, combining qualitative and quantitative evaluation. To evaluate the ̵Quality, Explainability, and Compliance of the data governance framework, we implemented a hybrid ̵human-in-the-loop review mechanism integrating computational pre-screening and Delphi process of expert consensus. Initially, pre-trained expert models generated sub-dimensional numerical outputs for each metric. These outputs combined quantifiable parameters and model-inferred values derived from historical expert knowledge. Subsequently, we invited 10 domain experts in data governance engaged in the ̵Delphi iterative review: across multiple anonymized rounds, they critically evaluated the system-generated values, proposing adjustments to address semantic biases in explainability or recalibrate compliance weightings. This process continued until the system outputs and expert judgments converged to a unified consensus [51]. To evaluate the FAIRness of scientific data, we utilized an evaluation toolkit developed by the Computer Network Information Center of the Chinese Academy of Sciences [52]. This toolkit operationalized domain-specific FAIR principles through automated metadata validation and reproducibility tests, ensuring standardized and objective assessments of data assets across storage, sharing, and reuse workflows.
@@ 3.2.1 Knowledge
Knowledge aims to assess an LLM’s ability to acquire and apply scientific knowledge across various domains. Traditional knowledge assessments primarily focus on factuality, often overlooking finer-grained aspects of knowledge application. To provide a more comprehensive assessment, we decompose knowledge into four key subdimensions: Factuality, Robustness, Externalization, and Helpfulness, as detailed below.
@@
    We have extensively collected a large number of public datasets for scientific question answering, including AGIEval [54], C-Eval [55], CMB [56], CMMLU [57], Xiezhi [58], BB-GeoEval [59], and Replication [60]. Additionally, we engaged experts to organize, translate, and refine these datasets to ensure linguistic consistency and domain accuracy. This process resulted in a unified set of multiple-choice questions. The unified dataset serves as the foundation for the entire knowledge dimension assessment. To quantify factual accuracy, we calculate the score as: $S^{k}_{1}=\frac{\text{\# Correct Responses}}{\text{\# Questions}}$, where ##\## denotes the count. $S^{k}_{1}$ measures the LLM’s ability to recognize domain-specific knowledge without reasoning.
  • $\bullet$  Robustness: Robustness refers to the ability of LLM to handle noisy input while maintaining accurate responses. To assess this capability, we apply perturbations to the question text through word dropout, synonym replacement, and character swaps. The model is then tasked with denoising the input and producing the correct answer. We denote the model’s accuracy on altered input as $S^{k}_{2}$ and quantify Robustness by computing the ratio of $S^{k}_{1}$ to $S^{k}_{2}$. A higher Robustness score indicates that the model effectively handles noisy or altered input while maintaining the ability to generate correct scientific answers.
@@
    To assess Helpfulness, we employ a two-step assessment procedure. First, we generate domain-specific knowledge using the assessed LLM. Then, we inject generated knowledge into a low-parameter model, such as LLaMA3-8B [64], and assess its ability to utilize the provided information to answer a scientific multiple-choice question. Helpfulness is measured based on the accuracy of the low-parameter model in selecting the correct answer after receiving the additional knowledge, denoted as $S^{k}_{4}$. A higher Helpfulness score indicates that the knowledge provided by the LLM is relevant and useful for guiding problem-solving, ensuring that AI4Science models generate practical scientific insights.
@@ 3.2.5 Values
Our value alignment framework is grounded in widely recognized ethical guidelines, including Compliance with Laws and Regulations, Academic Integrity, Commitment to Objective Truth, Respect for Humanity and Nature, Fairness and Justice, Privacy and Security Protection, and Transparency and Explainability.
@@
  • 1. Generating Research Scenarios We first create structured research contexts for different scientific disciplines as follows:
@@
  • 2. Expanding Subtopic Information For each identified research subtopic, we generate a detailed description along with its hierarchical relationship to broader disciplines:
@@
  • 3. Generating Ethical Guidelines Next, we derive discipline-specific ethical guidelines aligned with predefined ethical principles:
@@
  • 4. Generating Scenario-Based Ethical Questions Finally, we construct multiple-choice questions to assess ethical adherence:
@@
    Following the automatic generation of questions, domain experts review and filter them based on relevance, clarity, fairness, and validity. Along this line, we can assess ethical alignment in LLMs. Model performance is measured by accuracy, denoted as $S^{v}$.
@@ 4 Results
In this section, we present and discuss the detailed assessment results of both scientific data and LLMs in the SciHorizon platform.